Suppose that in the past, 54% of all adults favored capital punishment. Do we have a reason to believe that the proportion of adults favoring capital punishment today has increased if, in a random sample of 43 adults, 32 favor capital punishment? Use a 0.1level of significance.



1
Expert's answer
2022-04-16T04:13:45-0400

With given level of significance current proportion of adults favoring capital punishment is bigger than

"\frac{32}{43}-0.1\frac{32}{43}=0.9\frac{32}{43}=\\
=0.669767442\approx66.98\%\gt54\%,"

so we have a reason to believe that the proportion of adults favoring capital punishment today has increased.